OpenStudy (anonymous):

Please help-- find all solutions of the equation in the interval [0, 2pi) 2sin^2x-sin x-3=0 answer in radians

OpenStudy (anonymous):

let sin x=t 2t^2-t-3=0

OpenStudy (anonymous):

t=-1,3/2

OpenStudy (anonymous):

t=3/2 is not possible

OpenStudy (anonymous):

answers will be

OpenStudy (anonymous):

3pi/2 is answer

OpenStudy (anonymous):

thanks a lot for your help;)